OpenStudy (anonymous):

Find the Domain and Range of y=2x/(x-1)

OpenStudy (anonymous):

your domain is anything that would make y undefined... which would mean that x-1=0

OpenStudy (anonymous):

the function becomes undifined when denominator is 0

OpenStudy (anonymous):

so x -1 not equal 0 x not equal 1 Domain = R-{1}

OpenStudy (anonymous):

The range is what I actually needed. Domain of fractions are pretty simple.

OpenStudy (anonymous):

the range is the y where x is in the domain

OpenStudy (anonymous):

for Range put y = f(x) y=2x/(x-1) y(x-1) = 2x yx - y = 2x x (y-2) = y x = y/y-2

OpenStudy (anonymous):

nw Same Condition Apply y - 2 not equal 0 y not equal 2 so Range = R-{2}

OpenStudy (anonymous):

@Grazes got it ?

OpenStudy (anonymous):

How did yx - y = 2x turn into x (y-2) = y

OpenStudy (anonymous):

yx - 2x = y x (y-2) = y

OpenStudy (anonymous):

range is everything actually....

OpenStudy (anonymous):

Ahh. Alright.

OpenStudy (anonymous):

cuz at the asymptote..... it goes up to positive and dwon to negative infinities

OpenStudy (anonymous):

with the horizontal one at 2

OpenStudy (anonymous):

sooo Range = (-infinity,2);(2,infinity)

OpenStudy (anonymous):

Range is {y|y∈ℝ, y≠2} According to Yahoo

OpenStudy (anonymous):

which is exactly what i stated... though in a different way... cuz basically what Yahoo is saying is that range is everything except for when y=2

OpenStudy (anonymous):

yepyepyep

OpenStudy (anonymous):

Could you also divide 2x by x-1 using polynomial division?